5 years ago,amit was 3 times old as armaan.10 years later amit would be twice old as armaan.how old are they now?

Answer by anand429(138)

let amit's current age = a and armaan's current age be b.
As per ques,
(a-5) = 3* (b-5)
i.e. a= 3b-10 -(i)
10 years later,
(a+10) = 2(b+10)
i.e. a=2b+10 -(ii)
From eqn (i) and (ii),
b=20,
and a=50